# Practical approaches for crystal structure predictions with inpainting generation and universal interatomic potentials

### v1

@@ Abstract
We present Crystal Host-Guided Generation (CHGGen), a diffusion-based framework for crystal structure prediction. Unconditional generation with diffusion models demonstrates limited efficacy in identifying symmetric crystals as the unit cell size increases. CHGGen addresses this limitation through conditional generation with the inpainting method, which optimizes a fraction of atomic positions within a predefined and symmetrized host structure. We demonstrate the method on the ZnS-P2S5 and Li-Si chemical systems, where the inpainting method generates a higher fraction of symmetric structures than unconditional generation. The practical significance of CHGGen extends to enabling the structural modification of crystal structures, particularly for systems with partial occupancy, surface absorption and defects. The inpainting method also allows for seamless integration with other generative models, providing a versatile framework for accelerating materials discovery.
@@ I Introduction
Recent advances in graph neural network (GNN)-based machine learning models have introduced promising alternatives to traditional CSP methods, with a key milestone being the development of universal machine learning interatomic potentials (uMLIPs)—offering accurate and transferable modeling across diverse material systems [11, 12, 13, 14]. These uMLIPs trained on millions of DFT calculations demonstrate remarkable generalizability in exploring vast chemical spaces for materials discovery [15, 16, 17, 18]. Another emerging direction is deep-generative models, particularly diffusion models, which learn the data manifold or probabilistic distribution and generate new configurations via stochastic or variational approaches [19, 20]. Xie et al. [21] introduced CDVAE that uses a variational autoencoder to sample lattice parameters and compositions and a diffusion model to optimize atomic coordinates. Although promising, CDVAE-generated structures are predominantly thermodynamically unstable or lack symmetry [22, 23]. Kurz et al. [24] introduced a wrapped normal distribution to effectively couple lattice diffusion with fractional coordinates, a strategy that was successfully implemented in DiffCSP. [24, 25, 26]. Zeni et al. [27] further adapted the scheme with edge features for lattice scores in the MatterGen framework that enables diffusions on lattices, fractional coordinates, and chemical species. By learning from materials datasets such as Materials Project (MP) [28] and Alexandria [29], MatterGen is capable of conducting scalable and universal exploration in high-dimensional design space and achieves excellent performance in structural stability, uniqueness, and novelty, albeit with an observed limitation to a smaller scale (e.g., $N_{\text{atom}}\leq 20$). Beyond GNN-based models, other approaches such as U-net-based diffusion models [30], optimization of subcell structures from amorphous configurations [31], and large language models [32, 33] have shown promise in crystal structure generations without the need to limit structure sizes.
@@
In this work, we extend GNN-based diffusion models to enable fractional crystal structure design via inpainting generation—that is, given a host or substrate structure, we optimize the placement of additional ’guest’ atoms within the existing framework. This application is particularly valuable in several material domains, e.g., defective materials, intercalation electrodes [34], molecular absorption on catalyst surfaces [35], and interfacial solid reactions where surfaces reconstruct while bulk structures remain unchanged [36]. We first summarize the fundamentals of diffusion models and inpainting generation, and discuss the locality bias of GNN-based diffusion models, particularly as a key challenge when generation is performed at large scales. To address these gaps, we introduce Crystal Host-Guided Generation (CHGGen), which integrates inpainting generation based on symmetrized frameworks and uMLIP for structure optimization. We demonstrate the effectiveness of host-guided generation through a case study on CSP within the ZnS–P2S5 chemical space, and showcase the broader applicability of CHGGen across the continuous chemical space of the Lix–Si alloy system. Finally, we discuss the limitations and potential opportunities of applying CHGGen with state-of-the-art generative models for CSP problems in future directions.
@@ II Theory
Diffusion model.
Generating samples from a probability density function $p(\boldsymbol{x})$ in high-dimensional space $\mathbb{R}^{d}$ can be achieved by modeling the gradient of the log-probability density, known as the score function $\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x})$ in diffusion models. Song et al. [37] demonstrated that both the diffusion process and its reverse can be formulated as stochastic differential equations (SDE)
@@
Denoising score matching.
To estimate the score function $\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})$, we use score matching (SM) to optimize the model parameters $\theta^{*}$ by minimizing
@@
Here $\boldsymbol{x}(0)\sim p_{0}(\boldsymbol{x})$ and $\boldsymbol{x}(t)\sim p(\boldsymbol{x}(t)|\boldsymbol{x}(0))$, $s_{\theta}(\boldsymbol{x}(t),t)$ is the score function predicted by the GNN model, $\boldsymbol{\epsilon}$ represents the normalized noise $[\boldsymbol{x}(t)-\boldsymbol{x}(0)]/\sigma\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. In the training process, $\sigma$ is sampled uniformly from the interval $[\sigma_{\min},\sigma_{\max}]$ to perturb the configuration $\boldsymbol{x}(0)$ and obtain the noisy configuration $\boldsymbol{x}(t)$ to construct the DSM loss in Eq. (6).
Inpainting.
Inpainting is a conditional generation process where a model completes missing elements within a given context. Originally developed for image processing [40], inpainting has demonstrated significant applications in materials and chemistry, including the discovery of chemical reaction transition states [41] and the generation of symmetry-constrained 2D materials [42]. In CSP, inpainting enables the optimal placement of additional atoms (termed guest atoms) within a predefined host crystal structure.
Lugmayr et al. [40] introduced the repaint algorithm for high-quality image inpainting using diffusion models. The approach leverages standard unconditional diffusion models trained with DSM. During inference, repaint employs two key mechanisms: (1) a masking strategy that applies different noise treatments to known regions (host structure) and unknown regions (areas to be inpainted – guest atoms), and (2) a resampling technique that repeatedly ”jumps back” in the diffusion process and resamples the unknown regions multiple times at each timestep $t$. This resampling procedure helps harmonize the generated content with existing regions by allowing multiple attempts at generating coherent inpainted content. For detailed implementation and theoretical foundations, we refer readers to Ref. [40] and [43] for details of this approach.
@@
This ensures that both the guest atoms and host structure in the current configuration $\boldsymbol{x}_{t}$ maintain comparable noise scales. The GNN model then computes the score using all atoms and executes the reverse diffusion via Eq. (4) (Process B). In our notation, $\boldsymbol{x}^{\mathrm{host}}$ contains atomic positions of all atoms (including the framework and guest atoms) for indexing consistency. We apply masks $(1-m)$ to $\boldsymbol{x}^{\mathrm{host}}$ and $m$ to $\boldsymbol{x}$ (Process C) to construct $\boldsymbol{x}_{t-1}$ (Process D):
@@
Through this iterative reverse diffusion process, the noise scale $\{\sigma_{t}\}$ gradually decreases, resulting in a final crystal structure that is closely aligned with the original host structure with minimal deviation (e.g., $\sigma_{\text{min}}=0.001$). The positions of the guest atoms are therefore determined by the distribution conditioned on the host structure.
@@
III Results
We developed an SE(3)-equivariant Graph Neural Network (GNN) based on the NequIP architecture to predict the score function for both unconditional and inpainting diffusion processes. The model training followed the DSM scheme described in Eq. (6), where the dataset was prepared using crystal structures from the MP database with energy above hull $E_{\text{hull}}<0.1$ eV (see Methods).
In the following sections, we first examine the locality bias encountered when generating structures with large unit cells through unconditional generation. These insights led to the development of the CHGGen framework. We demonstrate the effectiveness of CHGGen on two example chemical systems: Zn-P-S and Li-Si, which are complemented by CHGNet as a uMLIP for iterative structural optimizations and thermodynamic stability screening.
@@ III.1 Locality bias of GNN-based generative models
To investigate the limitations of unconditional generation at large scales, we generated 10 supercell configurations of (Li2S-P2S5)10 and (Li4S2-P2S5)8 with the parametrized diffusion models (Fig. 2a). Figure 2a shows a snapshot of the generated structures, where the long-range periodicity is absent and an amorphous configuration is exhibited. Figure  2c and 2d present the radial distribution functions (RDF) of the generated structures and the structures from the MP database, showing similar major peaks in both RDF plots. The RDF suggests that the generated structures exhibit physically reasonable Li–S and P–S bonding environments, which represent the learned local distribution from the dataset.
@@ III.2 Crystal host-guided generation
Figure 4: Generation results in the Zn-P-S system demonstrating local stability, global stability, and capability for identifying symmetric structures. (a) Energy change ($\Delta E$) following structure relaxation. (b) Maximum pair-wise root-mean-squared displacement (RMSD) representing differences between generated and relaxed structures. (c) Distribution of decomposition energies ($E_{d}$) of generated structures relative to the MP phase diagram as predicted by CHGNet. (d–e) Comparison of decomposition energies predicted by CHGNet (blue) and r2SCAN-DFT (orange) relative to the MP phase diagram at GGA/GGA$+U$ (blue) and r2SCAN (orange) levels of accuracy. (f) Success rate of identifying symmetric crystal structures using unconditional vs. inpainting generation. (g) Examples of generated structures with the lowest DFT decomposition energy in ZnP2S6 and Zn2P2S7 compositions.
@@
Motivated by our observation that GNN-based diffusion models struggle to generate crystalline structures with long-range periodic structures, we developed the CHGGen framework as a targeted approach to mitigate this limitation. CHGGen integrates three key components: (1) unconditional structure generation, (2) inpainting generation based on symmetry-refined host structure, and (3) structural optimization using the CHGNet uMLIP.
@@
Figure 3 illustrates the CHGGen computational workflow. The process begins with sampling various Bravais lattices at a fixed volume through a random search over lattice constant ratios and angles (see Methods). The unit cell volume is determined as $N*V_{0}$, where $N$ represents the number of atoms and $V_{0}$ denotes the atomic volume. The $V_{0}$ can be initialized either from related crystalline phases or predicted by composition-based regression models [47]. This atomic volume serves as prior information subject to optimization in subsequent steps. Following lattice determination, fractional coordinates for all atoms are initialized with random numbers drawn from $\mathcal{N}(0,1)$. The diffusion process then proceeds by solving the reverse SDE using scores predicted by the SE(3)-GNN (unconditional generation). Given that the volumes and lattices of the generated structures are drawn from simple priors and random search, the CHGNet uMLIP is employed for structure relaxation to optimize both unit cells and atomic coordinates. This process represents a well-defined local energy minima search task and does not suffer from the locality bias encountered during the diffusion process.
@@
The next phase is initiated by removing atoms that exhibit broad local environment distributions (e.g., Li). The remaining structure (framework) undergoes symmetry refinement using spglib through incremental structural matching tolerance to obtain a space group with higher symmetry (i.e., until the space group is not P1). Since the guest atoms exhibit diverse local environment distributions, refinement without them is more feasible for obtaining the symmetric structure. The fractional coordinates of the removed guest atoms are then reinitialized from $\mathcal{N}(\mathbf{0},\mathbf{I})$ within the symmetrized framework, and inpainting generation is performed using masks $m$ and $(1-m)$ for the guest and framework atoms, respectively.
@@ III.3 Example: Zn-P-S
We focused on the CSP of ZnP2S6 and Zn2P2S7 using CHGGen. To assess the local stability of the generated structures, we evaluated the structural and energetic differences between the initially generated structures and their CHGNet-relaxed counterparts. In Fig. 4a and 4b, we present the energy and geometrical differences between the relaxed structures and generated structures. Most of the generated structures exhibit energy changes of $\Delta E<0.1$ eV/atom, with a median value of 0.08 eV/atom. The geometric differences, quantified by maximum pair-wise root-mean-squared distance (RMSD), show a median value of 0.10 Å between relaxed and generated structures. As illustrated in Fig. 4a-b, outliers with large energy changes correspond to RMSD values exceeding 0.3 Å, indicating that most of the generated structures are close to the local minima and can be reasonably searched using uMLIP structure relaxation.
@@
Figure 4c presents the CHGNet-predicted decomposition energies ($E_{d}$) to quantify the thermodynamic stability. The distribution shows a median $E_{d}$ of 0.07 eV/atom, indicating that the majority of generated structures are metastable in this chemical space, while 6.5% of the generated structures are identified as stable with $E_{d}<0$ at the uMLIP level of accuracy.
@@
After r2SCAN-DFT calculations and reevaluation of $E_{d}$ with respect to the MP-r2SCAN phase diagram, the structure with negative $E_{d}$ was confirmed to be metastable. This discrepancy can be attributed to CHGNet’s approximate 30 meV/atom prediction error, which may alter the predicted phase stability when $E_{d}$ values are small. As illustrated in Fig. 4d and 4e, while CHGNet demonstrates good overall agreement with DFT, it systematically underestimates the decomposition energy for structures with high values ($E_{d}>0.1$ eV/atom), consistent with the known softening effect of uMLIPs [49]. This finding suggests that practitioners could consider using a lower threshold (e.g., $E_{d}<30$ meV/atom) when screening generated structures for DFT validation.
@@
To evaluate how diffusion models perform in generating symmetric crystal structures, we define a simple metric symmetry success rate as the fraction of generated structures possessing space groups with higher symmetry than P1 or P-1 after relaxation and refinement (see Methods). Figure 4f compares the symmetry success rates from unconditional (blue) and inpainting (orange) generations. The inpainting approach uses the refined P–S frameworks to achieve significantly higher success rates compared to the unconditional ones ($<5\%$), which highlights the effectiveness and practical advantages of conditional generation with structural priors. The structures with the lowest-$E_{d}$ for ZnP2S6 and Zn2P2S7 are illustrated in Fig. 4g. While the generated structures are predicted to be metastable by DFT calculations, this example demonstrates CHGGen’s potential for exploring crystal structures in the chemical space that is currently absent from existing databases.

### v2

# Crystal structure prediction with host-guided inpainting generation and foundation potentials

@@ Abstract
Unconditional crystal structure generation with diffusion models faces challenges in identifying symmetric crystals as the unit cell size increases. We present the Crystal Host-Guided Generation (CHGGen) framework to address this challenge through conditional generation using an inpainting method, which optimizes a fraction of atomic positions within a predefined and symmetrized host structure to improve the success rate for symmetric structure generation. By integrating inpainting structure generation with a foundation potential for structure optimization, we demonstrate the method on the ZnS-P2S5 and Li-Si chemical systems, where the inpainting method generates a higher fraction of symmetric structures than unconditional generation. The practical significance of CHGGen extends to enabling the structural modification of crystal structures, particularly for systems with partial occupancy or intercalation chemistry. The inpainting method also allows for seamless integration with other generative models, providing a versatile framework for accelerating materials discovery.
@@ I Introduction
Recent advances in graph neural network (GNN)-based machine learning models have introduced promising alternatives to traditional CSP methods, with a key milestone being the development of foundation potential, or universal machine learning interatomic potentials—offering accurate and transferable modeling across diverse material systems [11, 12, 13, 14]. These foundation potentials trained on millions of DFT calculations demonstrate remarkable generalizability in exploring vast chemical spaces for materials discovery [15, 16, 17, 18]. Another emerging direction is deep-generative models, particularly diffusion models, which learn the data manifold or probabilistic distribution and generate new configurations via stochastic or variational approaches [19, 20, 21]. Xie et al. [22] introduced CDVAE that uses a variational autoencoder to sample lattice parameters and compositions and a diffusion model to optimize atomic coordinates. Although promising, CDVAE-generated structures are predominantly thermodynamically unstable or lack symmetry [23, 24]. Kurz et al. [25] introduced a wrapped normal distribution to effectively couple lattice diffusion with fractional coordinates, a strategy that was successfully implemented in DiffCSP. [25, 26, 27]. Zeni et al. [28] further adapted the scheme with edge features for lattice scores in the MatterGen framework that enables diffusions on lattices, fractional coordinates, and chemical species. By learning from materials datasets such as Materials Project (MP) [29] and Alexandria [30], MatterGen is capable of conducting scalable and universal exploration in high-dimensional design space and achieves excellent performance in structural stability, uniqueness, and novelty, albeit with an observed limitation to a smaller scale (e.g., $N_{\text{atom}}\leq 20$). Beyond GNN-based models, other approaches such as U-net-based diffusion models [31], optimization of subcell structures from amorphous configurations [32], and large language models [33, 34] have shown promise in crystal structure generations without the need to limit structure sizes.
@@
In this work, we extend GNN-based diffusion models to enable fractional crystal structure design via inpainting generation—that is, given a host or substrate structure, we optimize the placement of additional ’guest’ atoms within the existing framework. This application is particularly valuable in several material domains, e.g., defective materials, intercalation electrodes [35], molecular absorption on catalyst surfaces [36], and interfacial solid reactions where surfaces reconstruct while bulk structures remain unchanged [37]. We first summarize the fundamentals of diffusion models and inpainting generation, and discuss the locality bias of GNN-based diffusion models, particularly as a key challenge when generation is performed at large scales. To address these gaps, we introduce Crystal Host-Guided Generation (CHGGen), which integrates inpainting generation based on symmetrized frameworks and foundation potential for structure optimization. We demonstrate the effectiveness of host-guided generation through a case study on CSP within the ZnS–P2S5 chemical space, and showcase the broader applicability of CHGGen across the continuous chemical space of the Lix–Si alloy system. Finally, we discuss the limitations and potential opportunities of applying CHGGen with state-of-the-art generative models for CSP problems in future directions.
@@ II Theory
Diffusion model. Generating samples from a probability density function $p(\boldsymbol{x})$ in high-dimensional space $\mathbb{R}^{d}$ can be achieved by modeling the gradient of the log-probability density, known as the score function $\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x})$ in diffusion models. Song et al. [38] demonstrated that both the diffusion process and its reverse can be formulated as stochastic differential equations (SDE)
@@
Denoising score matching. To estimate the score function $\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})$, we use score matching (SM) to optimize the model parameters $\theta^{*}$ by minimizing
@@
Here $\boldsymbol{x}(0)\sim p_{0}(\boldsymbol{x})$ and $\boldsymbol{x}(t)\sim p(\boldsymbol{x}(t)|\boldsymbol{x}(0))$, $s_{\theta}(\boldsymbol{x}(t),t)$ is the score function predicted by the GNN model, $\boldsymbol{\epsilon}$ represents the normalized noise $\boldsymbol{x}(t)-\boldsymbol{x}(0)]/\sigma\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. In the training process, $\sigma$ is sampled uniformly from the interval $\sigma_{\min},\sigma_{\max}]$ to perturb the configuration $\boldsymbol{x}(0)$ and obtain the noisy configuration $\boldsymbol{x}(t)$ to construct the DSM loss in Eq. (6).
Inpainting. Inpainting is a conditional generation process where a model completes missing elements within a given context. Inpainting has demonstrated significant applications in materials and chemistry, including the discovery of chemical reaction transition states [41] and the generation of symmetry-constrained 2D materials [42]. In CSP, inpainting enables the optimal placement of additional atoms (termed guest atoms) within a predefined host crystal structure, where a binary masking strategy applies different noise treatments to known regions (host structure) and unknown regions (areas to be inpainted – guest atoms).
Unlike training a certain distribution of the mask, Lugmayr et al. [43] introduced the repaint algorithm (inpainting + resampling) for high-quality 2D image inpainting using diffusion models. One can simply train the diffusion model with DSM to learn the joint distribution. During inference, the conditional distribution is approached using the resampling technique for inpainting generation. As shown in the Algorithm, in addition to the unconditional generation steps, the resampling repeatedly ”jumps back” in the diffusion process and resamples the unknown regions multiple times with $r$ steps) at each timestep $t$, with a mask $m$ to separate the host and guest atoms. This resampling procedure helps harmonize the generated content with existing regions by allowing multiple attempts at generating coherent inpainted content. For detailed implementation and theoretical foundations, we refer readers to Ref. [43] and [44] for details of this approach.
@@
This ensures that both the guest atoms and host structure in the current configuration $\boldsymbol{x}_{t}$ maintain comparable noise scales. The GNN model then computes the score using all atoms and executes the reverse diffusion via Eq. (4) (Process B). In our notation, $\boldsymbol{x}^{\mathrm{host}}$ contains atomic positions of all atoms (including the framework and guest atoms) for indexing consistency. We apply masks (1m)(1-m) to $\boldsymbol{x}^{\mathrm{host}}$ and $m$ to $\boldsymbol{x}$ (Process C) to construct $\boldsymbol{x}_{t-1}$ (Process D):
@@
Through this iterative reverse diffusion process, the noise scale $\{\sigma_{t}\}$ gradually decreases, resulting in a final crystal structure that is closely aligned with the original host structure with minimal deviation (e.g., $\sigma_{\text{min}}=0.001$ Å). The positions of the guest atoms are therefore determined by the distribution conditioned on the host structure.
III Results
We developed an SE(3)-equivariant Graph Neural Network (GNN) based on the NequIP architecture to predict the score function for both unconditional and inpainting diffusion processes. The model training followed the DSM scheme described in Eq. (6), where the dataset was prepared using crystal structures from the MP database with energy above hull $E_{\text{hull}}<0.1$ eV (see Methods in SI).
In the following sections, we first examine the locality bias encountered when generating structures with large unit cells through unconditional generation. These insights led to the development of the CHGGen framework. We demonstrate the effectiveness of CHGGen on two example chemical systems: Zn-P-S and Li-Si, which are complemented by CHGNet as a foundation potential for iterative structural optimizations and thermodynamic stability screening. Additionally, we also demonstrate example studies of CSP of 16 compositions used in DiffCSP [27] and solid-solid interface in Supplementary Information.
@@ III.1 Locality bias of GNN-based generative models
To investigate the limitations of unconditional generation at large scales, we generated 10 supercell configurations of (Li2S-P2S5)10 and (Li4S2-P2S5)8 with the parametrized diffusion models (Fig. 2a). Figure 2a shows a snapshot of the generated structures, where the long-range periodicity is absent and an amorphous configuration is exhibited. Figure 2c and 2d present the radial distribution functions (RDF) of the generated structures and the structures from the MP database, showing similar major peaks in both RDF plots. The RDF suggests that the generated structures exhibit physically reasonable Li–S and P–S bonding environments, which represent the learned local distribution from the dataset.
@@ III.2 Crystal host-guided generation
Figure 4: Generation results in the Zn-P-S system demonstrating local stability, global stability, and capability for identifying symmetric structures. (a) Energy change ($\Delta E$) following structure relaxation. (b) Maximum pair-wise root-mean-squared displacement (RMSD) representing differences between generated and relaxed structures. (c) Distribution of decomposition energies ($E_{d}$) of generated structures relative to the MP phase diagram as predicted by CHGNet. (d–e) Comparison of decomposition energies predicted by CHGNet (blue) and r2SCAN-DFT (orange) relative to the MP phase diagram at GGA/GGA+U+U (blue) and r2SCAN (orange) levels of accuracy. (f) Success rate of identifying symmetric crystal structures using unconditional vs. inpainting generation. (g) Examples of generated structures with the lowest DFT decomposition energy in ZnP2S6 (C2, $E_{d}$ = 0.015 eV/atom) and Zn2P2S7 (Cm, $E_{d}$ = 0.046 eV/atom)
@@
Motivated by our observation that GNN-based diffusion models struggle to generate crystalline structures with long-range periodic structures, we developed the CHGGen framework as a targeted approach to mitigate this limitation. CHGGen integrates three key components: (1) unconditional structure generation, (2) inpainting generation based on symmetry-refined host structure, and (3) structural optimization using the CHGNet.
@@
Figure 3 illustrates the CHGGen computational workflow. The process begins with sampling various Bravais lattices at a fixed volume through a random search over lattice constant ratios and angles (see Methods in SI). The unit cell volume is determined as $N*V_{0}$, where $N$ represents the number of atoms and $V_{0}$ denotes the atomic volume. The $V_{0}$ can be initialized either from related crystalline phases or predicted by composition-based regression models [48]. This atomic volume serves as prior information subject to optimization in subsequent steps. Following lattice determination, fractional coordinates for all atoms are initialized with random numbers drawn from $\mathcal{N}(0,1)$. The diffusion process then proceeds by solving the reverse SDE using scores predicted by the SE(3)-GNN (unconditional generation). Given that the volumes and lattices of the generated structures are drawn from simple priors and random search, the CHGNet is employed for structure relaxation to optimize both unit cells and atomic coordinates. This process represents a well-defined local energy minima search task and does not suffer from the locality bias encountered during the diffusion process.
@@
The next phase is initiated by removing atoms that exhibit broad local environment distributions (e.g., Li). The remaining structure (framework) undergoes symmetry refinement using spglib through incremental structural matching tolerance to obtain a space group with higher symmetry (i.e., until the space group is not P1). Since the guest atoms exhibit diverse local environment distributions, refinement without them is more feasible for obtaining the symmetric structure. The fractional coordinates of the removed guest atoms are then reinitialized from $\mathcal{N}(\mathbf{0},\mathbf{I})$ within the symmetrized framework, and inpainting generation is performed using masks $m$ and (1m)(1-m) for the guest and framework atoms, respectively.
@@ III.3 Example: Zn-P-S
We focused on the CSP of ZnP2S6 and Zn2P2S7 using CHGGen. To assess the local stability of the generated structures, we evaluated the structural and energetic differences between the initially generated structures and their CHGNet-relaxed counterparts. In Fig. 4a and 4b, we present the energy and geometrical differences between the relaxed structures and generated structures. Most of the generated structures exhibit energy changes of $\Delta E<0.1$ eV/atom, with a median value of 0.08 eV/atom. The geometric differences, quantified by maximum pair-wise root-mean-squared distance (RMSD), show a median value of 0.10 Å between relaxed and generated structures. As illustrated in Fig. 4a-b, outliers with large energy changes correspond to RMSD values exceeding 0.3 Å, indicating that most of the generated structures are close to the local minima and can be reasonably searched using foundation potential structure relaxation.
@@
Figure 4c presents the CHGNet-predicted decomposition energies ($E_{d}$) to quantify the thermodynamic stability. The distribution shows a median $E_{d}$ of 0.07 eV/atom, indicating that the majority of generated structures are metastable in this chemical space, while 6.5% of the generated structures are identified as stable with $E_{d}<0$ at the CHGNet level of accuracy.
@@
After r2SCAN-DFT calculations and reevaluation of $E_{d}$ with respect to the MP-r2SCAN phase diagram, the structure with negative $E_{d}$ was confirmed to be metastable. This discrepancy can be attributed to CHGNet’s approximate 30 meV/atom prediction error, which may alter the predicted phase stability when $E_{d}$ values are small. As illustrated in Fig. 4d and 4e, while CHGNet demonstrates good overall agreement with DFT, it systematically underestimates the decomposition energy for structures with high values ($E_{d}>0.1$ eV/atom), consistent with the known softening effect of foundation potentials [50]. This finding suggests that practitioners could consider using a lower threshold (e.g., $E_{d}<30$ meV/atom) when screening generated structures for DFT validation.
@@
To evaluate how diffusion models perform in generating symmetric crystal structures, we define a simple metric symmetry success rate as the fraction of generated structures possessing space groups with higher symmetry than P1 or P-1 after relaxation and refinement (see Methods in SI). Figure 4f compares the symmetry success rates from unconditional (blue) and inpainting (orange) generations. The inpainting approach uses the refined P–S frameworks to achieve significantly higher success rates compared to the unconditional ones ($5\%$), which highlights the effectiveness and practical advantages of conditional generation with structural priors. The structures with the lowest-$E_{d}$ for ZnP2S6 and Zn2P2S7 are illustrated in Fig. 4g. While the generated structures are predicted to be metastable by DFT calculations, this example demonstrates CHGGen’s potential for exploring crystal structures in the chemical space that is currently absent from existing databases.